OpenStudy (lina777):

Please help! I have been trying to work this out for an hour and I am going crazy. I will award and medal!!! If sinxcos(pi/7)-sin(pi/7)cosx=-(sqrt2)/2, then x can equal __; check all that apply A. 3pi/4+pi/7+2npi B. 7pi/4+pi/7+2npi C. pi/4+pi/7+2npi D. 5pi/4+pi/7+2npi

jimthompson5910 (jim_thompson5910):

look at this list of trig identities http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf look on page 2 at the section `Sum and Difference Formulas`

jimthompson5910 (jim_thompson5910):

which of those identities in that section will we use?

jimthompson5910 (jim_thompson5910):

any ideas @lina777 ?

OpenStudy (lina777):

Thank you for answering! @jim_thompson5910 The identity we will use is sin(a-b)=sinacosb+cosasinb

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

it's actually sin(A-B) = sin(A)*cos(B) - cos(A)*sin(B) you will have a minus on the right side. Not a plus

jimthompson5910 (jim_thompson5910):

that means `sin(x)*cos(pi/7) - cos(x)*sin(pi/7)` will turn into `sin(x-pi/7)`

jimthompson5910 (jim_thompson5910):

so sin(x-pi/7) = -sqrt(2)/2 are you able to see what comes next?

jimthompson5910 (jim_thompson5910):

hint: it will involve the unit circle

OpenStudy (lina777):

Hello @jim_thompson5910 I got the stomach flu, I am sorry for not responding. Can you still help me out?

jimthompson5910 (jim_thompson5910):

so we left off at `sin(x-pi/7) = -sqrt(2)/2` right?

OpenStudy (lina777):

Right. :)

jimthompson5910 (jim_thompson5910):

Now use the unit circle https://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/2000px-Unit_circle_angles_color.svg.png to figure out which points have a y coordinate of \(\Large \frac{-\sqrt{2}}{2}\) Hint: there are 2 of these points. One is in Q3 and the other is in Q4

OpenStudy (lina777):

7pi/4 and 5pi/4

jimthompson5910 (jim_thompson5910):

let's just focus on the 7pi/4 for now

OpenStudy (lina777):

So is that the answer? 7pi/4+pi/7+2npi and 5pi/4+pi/7+2npi

jimthompson5910 (jim_thompson5910):

sin(x-pi/7) = -sqrt(2)/2 will mean that x-pi/7 = 7pi/4+2n*pi the `+2n*pi` part ensures that we capture all of the coterminal angles. The n represents any integer (whole number).

jimthompson5910 (jim_thompson5910):

`So is that the answer? 7pi/4+pi/7+2npi and 5pi/4+pi/7+2npi ` yes

OpenStudy (lina777):

Wow, that was much simpler than I thought it was! Thank you very much. And that is for sure?

jimthompson5910 (jim_thompson5910):

`And that is for sure? ` yes, I used a graphing calculator to confirm

OpenStudy (lina777):

Thank you so much for your help!!

jimthompson5910 (jim_thompson5910):

no problem